OpenStudy (anonymous):

2.Ismael is comparing cell phone plans before upgrading his phone. Ameri-Mobile offers a low activation fee, but a high monthly payment. Cell-U-Later offers a lower monthly rate, but the activation fee is higher. Create a possible algebraic expression for both Ameri-Mobile and Cell-U-Later that shows the amount paid after an unknown amount of months have passed. Justify how you created those expressions and identify what each term and factor represents in terms of the cell phone plans

OpenStudy (anonymous):

@XxSweet_NothingsxX

OpenStudy (anonymous):

1st month (m=1) : 10+2 = 12 2nd month (m=2) : 10+2+2 = 14 3rd month (m=3) : 10+2+2+2 = 16 In this example I just selected a startfee of 10 USD, but if there's no figure given, you can replace this number by 'S' and forget about an actual value. Same applies to the monthly fee, which you can replace by F if you don't know a value.

OpenStudy (anonymous):

i hope that helps XD